three consecutive number even numbers add up to 234. find the three numbers
Answer by stanbon(75887)

Even numbers are all multiples of 2.
---------
1st even, 2x
2nd even, 2x+2
3rd even, 2x+4
-----------------
EQUATION:
2x + 2x+2 + 2x+4 = 234
6x + 6 = 234
6x = 228
x = 38
2x= 76 (1st even)
2x+2 = 78 (2nd even)
2x+4 = 80 (3rd even)
=========================
Cheers,
Stan H.